With one method of the procedure called acceptance sampling, a sample of items is randomly selected without replacement , and the entire batch is rejected if there is at least one defect. The Newport Gauge Company has just manufactured a batch of aircraft altimeters and 3% are defective.
Question: If the batch contains 4000 altimeters and 100 of them are selected without replacement and tested, what is the probability that the entire batch will be rejected?

Answer by stanbon(75887)

With one method of the procedure called acceptance sampling, a sample of items is randomly selected without replacement, and the entire batch is rejected if there is at least one defect.
------------------------
The Newport Gauge Company has just manufactured a batch of aircraft altimeters and 3% are defective.
Question: If the batch contains 4000 altimeters and 100 of them are selected without replacement and tested, what is the probability that the entire batch will be rejected?
---
P(reject) = P(>=1 defect in 100)
= 1 - P(no defects in 100)
= 1 - (0.97)^100
= 1 - 0.0476
= 0.9524
===============
Cheers,
Stan H.
===============